A fluid flow is given by : V = 10x^3 i - 8x^3 yj. Find the shear strain rate and state whether the flow is rotational or irrotationsl.
1
Expert's answer
2020-12-10T13:23:34-0500
"V=10x^3 \hat{i}-8x^3y\hat{j}"

Gradient of velocity which is a second rank tensor is

"L^T=\vec\nabla \vec V=\begin{bmatrix}
 \cfrac{dV_x}{dx} & \cfrac{dV_x}{dy} \\
 \cfrac{dV_y}{dx} & \cfrac{dV_y}{dy}
\end{bmatrix}"

"L^T=\begin{bmatrix}
 30x^2 & 0 \\
 -24x^2y & -8x^3
\end{bmatrix}"

Taking Its transpose

"L=\begin{bmatrix}
 30x^2 & -24x^2y \\
 0 & -8x^3
\end{bmatrix}"

Now strain rate tensor:

"E=\cfrac{L+L^T}{2}\\
E=\cfrac{1}{2}\begin{bmatrix}
 60x^2 & -24x^2y \\
 -24x^2y & -16x^3
\end{bmatrix}"



where, off-diagonal terms represent shear strain rate.


Now Spin Tensor:

"W=\cfrac{L-L^T}{2}\\
W=\cfrac{1}{2}\begin{bmatrix}
 0 & -24x^2y \\
 24x^2y & 0
\end{bmatrix}"

Since off-diagonal term in spin tensor are non zero hence flow is rotational.